OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 10 when x = 4, write the direct linear variation equation. A. f(x) = 10x B. f(x) = 4x C. f(x) = 5/2x D. f(x) = 2/5x

OpenStudy (anonymous):

A. f(x) = 10x B. f(x) = 4x C. f(x) = 5/2x D. f(x) = 2/5x